Lactic acid, HC3H5O3 , is a weak acid in aqueous solution, Ka = 1.4 × 10−4 .

Calculate Kb for the lactate ion, C3H5O –
Answer #1

The product of Ka and Kb gives us the dissociation constant for water which is 1 x 10-7.

⇒ Ka x Kb = Kw = 1 x 10-7

Here, Ka = 1.4 × 10−4

⇒ Kb = 1 x 10-7 / 1.4 × 10−4 = 0.7143 x 10-3 = 7.143 x 10-4